OpenStudy (anonymous):

Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this. 9x – 2y = 88 -9x + 2y = -88

OpenStudy (hoblos):

if you multiply any one of the equations by (-1) you will get the second

OpenStudy (hoblos):

so you have an infinite number of solutions

OpenStudy (mathmagician):

multiply (-9x+2y=-88)by -1 and you get 9x-2y=88. So, there are infinite many solutions- it is any pair of (x;y) that statisfies equation 9x-2y=88

OpenStudy (anonymous):

So there is an infinite number of solutions and not just one specific answer to post?